DivineSolar (divinesolar):

Help will medal! Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y less than or greater to -x - 4 (1, 10) (-1, 10) (10, 1) (1, -10)

DivineSolar (divinesolar):

@sammixboo @mathstudent55 @phi

DivineSolar (divinesolar):

@CGGURUMANJUNATH

OpenStudy (phi):

You have to fix this " less than or greater to" things are not both less than and bigger than a number (example, you can't be both under 5ft tall and over 5 ft tall. It is one or the other) maybe you mean less than or equal ? or just less than ? or greater than or equal (written \( \geq \) ) or "greater than" written >

DivineSolar (divinesolar):

Equal to*

DivineSolar (divinesolar):

Yes less than or equal to

OpenStudy (phi):

<= or \( \leq \)

DivineSolar (divinesolar):

second choice ≤

DivineSolar (divinesolar):

refresh the page if u just see question marks.

OpenStudy (phi):

oh, I meant they mean the same thing. <= is how to type it if you can't type the fancy way. anyway, the system is : y \( \leq \) x - 5 y \( \leq \) -x - 4

DivineSolar (divinesolar):

Yes thats the system and the 4 answer choices are listed below.

OpenStudy (phi):

The easiest way is test the choices. Because we want y to be small, I would first test the choices with the smallest y value... hopefully that is the right answer.

DivineSolar (divinesolar):

C then?

OpenStudy (phi):

if y were temperature, is 1 degree above zero colder than -10 (10 below ) ?

DivineSolar (divinesolar):

No.

OpenStudy (anonymous):

phi could come to my Queation please

DivineSolar (divinesolar):

@,nmfrqew You dont just invade someones post to ask that.

OpenStudy (phi):

in other words, -10 is the smallest y but if you did test C we would use x=10 in the first equation y <= x - 5 y <= 10-5 ? y <= 5 y is 1 . it worked in the first relation. Now the 2nd: y <= -x - 4 replace x with 10 y <= -10 - 4 y <= -14 y is 1 for choice C (remember we are checking (10,1) ) is 1 <= -14. ?

DivineSolar (divinesolar):

no

OpenStudy (phi):

so I would test choice D. chances are it works.

OpenStudy (anonymous):

omg

DivineSolar (divinesolar):

the answer to your question nm is 0.5

DivineSolar (divinesolar):

Ok phi can you help with 1 moer?

DivineSolar (divinesolar):

more*

OpenStudy (anonymous):

srry if i was rude to u

DivineSolar (divinesolar):

Its okay.

OpenStudy (anonymous):

where is phi